Fourty-four percent of a corporation’s blue-collar employees

were in favor of a modified health care plan,

and 24% of its blue-collar employees favored a proposal

to change the work schedule. Thirty percent of

those favoring the health care plan modification, favored

the work schedule change.

a. What is the probability that a randomly selected

blue-collar employee is in favor of both the modified

health care plan and the changed work schedule?

b. What is the probability that a randomly chosen

blue-collar employee is in favor of at least one of

the two changes?

c. What is the probability that a blue-collar employee

favoring the work schedule change also favors the

modified health care plan?


1
Expert's answer
2021-12-16T15:03:28-0500

Let "A" be the event that employees are in favor of a modified heath plan and "B" be the event that employees favored a proposal to change the work schedule.

Given "P(A)=0.44, P(B)=0.24, P(B|A)=0.30"

a)


"P(B|A)=\dfrac{P(A\cap B)}{P(A )}"

"=>P(A\cap B)=P(B|A)P(A)"

"=0.3(0.44)=0.132"

"P(A\cap B)=0.132"

b)


"P(A\cup B)=P(A)+P(B)-P(A\cap B)"

"=0.44+0.24-0.132=0.548"

"P(A\cup B)=0.548"

c)


"P(A|B)=\dfrac{P(A\cap B)}{P(B )}=\dfrac{0.132}{0.24}=0.55"

"P(A|B)=0.55"